# TextFusion: Unveiling the Power of Textual Semantics for Controllable Image Fusion

### 21 Dec 2023 v1

@@ Abstract
Advanced image fusion methods are devoted to generating the fusion results by aggregating the complementary information conveyed by the source images. However, the difference in the source-specific manifestation of the imaged scene content makes it difficult to design a robust and controllable fusion process. We argue that this issue can be alleviated with the help of higher-level semantics, conveyed by the text modality, which should enable us to generate fused images for different purposes, such as visualisation and downstream tasks, in a controllable way. This is achieved by exploiting a vision-and-language model to build a coarse-to-fine association mechanism between the text and image signals. With the guidance of the association maps, an affine fusion unit is embedded in the transformer network to fuse the text and vision modalities at the feature level. As another ingredient of this work, we propose the use of textual attention to adapt image quality assessment to the fusion task. To facilitate the implementation of the proposed text-guided fusion paradigm, and its adoption by the wider research community, we release a text-annotated image fusion dataset IVT. Extensive experiments demonstrate that our approach (TextFusion) consistently outperforms traditional appearance-based fusion methods. Our code and dataset will be publicly available on the project homepage.
Figure 1: The input vision modalities and two fusion results obtained by our TextFusion. As reflected in the metrics, each application scenario requires a different fusion scheme to achieve the best performance. (EN is the image quality metric of information entropy and mAP denotes the mean average detection precision.)
1 Introduction
@@
Figure 2: Existing learning-based image fusion methods and the proposed controllable image fusion paradigm. To generate satisfactory fusion results for a specific scenario, existing methods require expensive retraining. The same goal can be achieved by simply adjusting the textual description in our paradigm.
The effectiveness of an image fusion approach will depend on the application scenes, i.e., whether it relates to, e.g., observation, detection, or recognition. As shown in Fig. 1, when image fusion is used in a detection task, the result (a) obtained by our method might be favoured, since it takes advantage of certain thermal radiation information to enhance the image contrast. On the other hand, result (b) is probably more appropriate for human observation, as it conveys more information (higher EN) and the image is rich in texture details (yellow boxes). However, such dependencies have not been studied in the existing literature. Instead, by relying purely on visual information, once trained, the current approaches are capable of producing fused images only in a fixed manner. Consequently, retraining fusion systems is required for different scenarios (Fig. 2 (a)), which limits their practical utility.
In this paper, we address the aforementioned lack of controllability. Specifically, our approach is to provide control over the fusion process using a textual description for the input vision modalities, which focuses on the fusion objectives. The text modality contains higher-level semantics for the image fusion task, overcoming the unwarranted reliance purely on the vision modalities. As shown in Fig. 2 (b), the proposed paradigm is able to deliver diverse image fusion outputs for different applications using a single image fusion model. A change of the fusion focus (concerned objects or regions) can be realised by simply adjusting the textual description for the input vision modalities. In general, our innovation introduces the use of linguistic information in the image fusion field.
@@
Figure 3: An illustration of the TextFusion model and the affine fusion unit designed to fuse the vision signals guided by the text modality.
2 Related work
2.1 Learning-based image fusion methods
@@
2.2 Vision-and-language pre-trained models
@@
2.3 Text prompt learning
@@ 3 Approach
3.1 Network architecture and the affine fusion unit
As shown in Fig. 3 (a), our network is composed of three encoders, i.e., two vision encoders and a text encoder. For the text encoder, we use the popular CLIP model to obtain linguistic features $\psi_{\mathrm{text}}$. On the other hand, the vision encoder part of CLIP contains down-sampling operations, which will impact the fusion performance when restoring the original size. Thus, we further design two vision encoders based on the Swin Transformer Blocks [25] to extract image features $\psi_{\mathrm{ir}}$ and $\psi_{\mathrm{vis}}$. After that, we insert an affine fusion unit into the data processing chain to integrate the extracted features. The final fused image is then reconstructed by a decoder, consisting of convolutional layers.
@@ 3.1 Network architecture and the affine fusion unit
Figure 4: Pipeline of the coarse-to-fine association mechanism.
3.2 Association and training processes
Coarse-to-fine association: As there is no GT image in the fusion task, to control the fused image via linguistic guidance, we need to embed the textual semantics into the loss function design. Thus, we need to generate an association map between the modalities for the training process. However, obtaining a pixel-wise relation map between vision and text modalities can be very expensive [9]. Thus, a coarse-to-fine association mechanism is proposed to alleviate this challenge.
As shown in Fig. 4, aiming to distinguish the user-defined content of interest in the input images from the background, we first construct a coarse-grained association (CGA) module to obtain the heat map $M$ linking each word to visual content, i.e., identifying the degree of correlation between the words and image regions. For each noun $\mathrm{w}_{j}$, we fuse their heat maps into a single one $\hat{M}(I)$ by taking the maximum value at each location of these heat maps, i.e.
@@ 3.2 Association and training processes
Taking this into consideration, we propose a subsequent fine-grained association (FGA) module to fine-tune the interest map. Such association is determined based on $Q$ segmentation instances of source images, e.g., persons, trees and cars. We first calculate the overlapping ratio $r$ between these instances and the interest map. Based on the threshold $\alpha$, we add the corresponding instance segmentation results together to generate the final interest map $B_{\mathrm{f}}$. Hence, this interest map is defined as:
@@
Training: During the training process, our loss function is designed to focus on the final interest map $B_{\mathrm{f}}$. The total loss is formulated as
@@
where $Loss_{\mathrm{int}}$ and $Loss_{\mathrm{irr}}$ stand for the region of interest and the irrelevant region loss components. As the thermal radiation information can be captured by the pixel intensity distribution, while the texture details in the RGB images are reflected in the gradient information [27, 46, 45], we use such criterion together with the principles of the dominant modality theory in the affine fusion unit to define these two loss functions. Specifically, the weights $w(B_{f},\phi(I))$ and $p(\overline{B_{f}},\nabla\theta(I))$ are adopted in the region of interest and irrelevant region, respectively, to measure the salience degree of source images. $\phi(I)$ and $\theta(I)$ denote the image features obtained from pre-trained models. $\nabla$ is the gradient operator and $\overline{B_{f}}(x,y)=1-B_{f}(x,y)$. For more details on the loss function, the reader is referred to the supplementary materials.
3.3 Textual attention assessments
@@
By utilizing the textual modality to describe the region of interest within the input images, we can embed a priori fusion knowledge into the IQAs to evaluate the fusion results. That is, depending on the purpose of the fusion task, we select salient information from the infrared modality and texture details from the visible images in the regions of interest and other areas, respectively, to construct a text-guided fused image $I_{\mathrm{tf}}$. We generate this image in a straightforward manner:
@@
4 Experimental results and analysis
4.1 Experimental setting
In this paper, our TextFusion is used to implement the infrared and visible image fusion task. As there is no available infrared and visible image dataset with the corresponding text modality, we carefully annotated data from several RGBT datasets, $i.e.$, TNO, RoadScene and LLVIP [35, 39, 11], to build a benchmark text-guided image fusion dataset (IVT). The final database which contains $11,405$ textual annotations will be released to the research community. More details about the annotation process and the participants (50 volunteers) are provided in the supplement.
During the training process, 10,000 RGBT and text triples from the IVT-LLVIP dataset were used to optimize the fusion model. The segmentation and the vision-and-language models are only pre-trained on the ImageNet classification tasks and the masked language modelling tasks [36, 13]. In the test phase, the entire IVT-TNO and IVT-LLVIP datasets, along with the remaining images from the IVT-LLVIP dataset, are used to evaluate the performance. For the quantitative comparison, some widely used metrics, i.e., an objective image fusion performance measure $Q_{abf}$, structural similarity (SSIM), and visual information fidelity (VIF) [42, 28] are chosen to measure the similarity of the fused image and source input from different perspectives. Besides, we also include two non-reference metrics, i.e., spatial frequency (SF) and standard deviation (SD) to evaluate the image quality of the fusion results. Larger values of these metrics indicate better fusion performance.
The comparison algorithms include the mainstream paradigms that emerged in the last few years, namely a GAN-based method, TarDAL [21], an autoencoder-based method, RFN-Nest [17], a self-evolution paradigm MUFusion [33], two methods combined with other vision tasks ReCoNet [10] and MetaFusion [21], a diffusion-based model DDFM [55] and a novel representation guided network, LRRNet [18]. The experiments are conducted on an NVIDIA GeForce RTX 3090 GPU and an Intel Core i7-6850K CPU.
@@ 4.2 Controllability characteristics
Table 1: The average detection precision of different fusion methods obtained on images from the IVT-LLVIP dataset. (Bold: best, Bold: second best)
@@
Figure 7: Sample qualitative results of our TextFusion compared with other state-of-the-art methods on the IVT dataset.
Table 2: Quantitative results of different image fusion methods obtained on the IVT dataset. ($mRank+$ and $mRank$ denote the mean rankings of the three textual attention and traditional metrics).
@@ 4.3 Pedestrian detection
Apart from the results visualisation, we also conducted quantitative experiments on this task. As presented in Table 1, with or without text input, our TextFusion always ranks first among these advanced image fusion methods.
@@
4.4 Image fusion results on the IVT dataset
@@
As shown in Table 2, compared to all advanced image fusion methods, our TextFusion delivers the best performance in terms of mRank on these three datasets. This demonstrates that, the results produced by our TextFusion maintain a high correlation with input images, and that our method is able to handle different fusion scenarios robustly. At the same time, as our method produces fused images with higher definition, it always achieves the best or second-best performance in terms of the no-reference metrics SF and SD. This objectively measured performance is consistent with the observations made in the qualitative evaluation.
@@
An analysis of the textual attention metrics: With our textual attention metrics, we aim to better measure the contributions of the source images to the fusion result quality controlled with the help of linguistic semantics. In this section, we combine the qualitative and quantitative experiments to demonstrate the effectiveness of the improved metrics. As shown in Fig. 7, the image fusion results of RFN-Nest suffer from the bias issue, i.e., the output either excessively preserves information from the RGB image or emphasizes prominent thermal radiation content at the expense of losing texture details. However, the traditional image fusion evaluation techniques, by evenly weighting the contributions from the source images, obscure this drawback. As a result, RFN-Nest still manages to achieve the second-best performance in terms of SSIM (Table 2). By contrast, our assessment (SSIM+) is able to identify this bias, with its average fusion ranking being suppressed to a certain extent (it drops from 6.00 to 6.67). Meanwhile, as MetaFusion and LRRNet can simultaneously preserve the important content from both modalities (though with lower quality than that produced by our method), their mRank benefits from the proposed assessments. In general, our textual attention metric can objectively identify the methods that fall short in certain aspects and promote those aligning with the fusion task objectives. This ensures that numerical results reflect more accurately the behaviour of image fusion approaches.
@@
Table 3: Quantitative results of the ablation study relating to the affine fusion unit.
@@
(c) $\mu$-vis $\lambda$-ir Moderate 0.491 0.657 0.637 17.897 45.175
Ours ($\mu$-ir $\lambda$-vis) Significant 0.527 0.656 0.744 18.872 49.578
4.5 Ablation study relating to the affine fusion unit

### 08 Feb 2024 v2

@@ Abstract
Advanced image fusion methods are devoted to generating the fusion results by aggregating the complementary information conveyed by the source images. However, the difference in the source-specific manifestation of the imaged scene content makes it difficult to design a robust and controllable fusion process. We argue that this issue can be alleviated with the help of higher-level semantics, conveyed by the text modality, which should enable us to generate fused images for different purposes, such as visualisation and downstream tasks, in a controllable way. This is achieved by exploiting a vision-and-language model to build a coarse-to-fine association mechanism between the text and image signals. With the guidance of the association maps, an affine fusion unit is embedded in the transformer network to fuse the text and vision modalities at the feature level. As another ingredient of this work, we propose the use of textual attention to adapt image quality assessment to the fusion task. To facilitate the implementation of the proposed text-guided fusion paradigm, and its adoption by the wider research community, we release a text-annotated image fusion dataset IVT. Extensive experiments demonstrate that our approach (TextFusion) consistently outperforms traditional appearance-based fusion methods. Our code and dataset will be publicly available at https://github.com/AWCXV/TextFusion.
Index Terms:
Image fusion, vision and language, multimodal learning, pre-trained model.
I Introduction
@@
Figure 1: The input modalities and two fusion results obtained by our TextFusion with different description. As reflected in the metrics, each application scenario requires a different fusion scheme to achieve the best performance. (EN is the image quality metric of information entropy and mAP denotes the mean average detection precision.)
The effectiveness of an image fusion approach will depend on the application scenes, i.e., whether it relates to, e.g., observation of the overall scene or area of interest, detection, or recognition. As shown in Fig. 1, when image fusion is used in a detection task, the result (a) obtained by our method might be favoured, since it takes advantage of certain thermal radiation information to enhance the image contrast. On the other hand, result (b) is probably more appropriate for human observation, as it conveys more information (higher EN) and the image is rich in texture details (yellow boxes). However, such dependencies have not been studied in the existing literature. Instead, by relying purely on visual information, once trained, the current approaches are capable of producing fused images only in a fixed manner. Consequently, retraining fusion systems is required for specific scenario or concerned objects (Fig. 2 (a)), which limits their practical utility.
Figure 2: Existing learning-based image fusion methods and the proposed controllable image fusion paradigm. To generate appropriate fusion results for a specific scenario (different tasks or concerned objects), existing methods cannot realise it or require expensive retraining. The same goal can be achieved by simply adjusting the focused objectives of textual description in our paradigm. (DET-all: general detection; OB-all: observation for the whole scene; OB-part: observation for the interested regions)
In this paper, we address the aforementioned lack of controllability by offering a variety of potential fusion style alternatives. Specifically, our approach is to provide control over the fusion process using a textual description for the input vision modalities, which focuses on the fusion objectives. The text modality contains higher-level semantics for the image fusion task, overcoming the unwarranted reliance purely on the vision modalities. As shown in Fig. 2 (b), the proposed paradigm is able to deliver diverse image fusion outputs for different scenes using a single image fusion model. A change of the fusion focus (concerned objects or regions) can be realised by simply adjusting the textual description for the input vision modalities. In general, our innovation introduces the use of linguistic information in the image fusion field.
@@
II Related work
II-A Learning-based image fusion methods
@@
II-B Vision-and-language pre-trained models
@@
II-C Text prompt learning
@@ IV Approach
Figure 5: An illustration of the TextFusion model and the affine fusion unit design. (a): Our fusion model receives the input image pairs and the textual description as input. The text encoder from the CLIP model and two vision encoders based on the Swin Transformer Blocks are used to extract the hidden representations of source input. We further propose an affine fusion model to align and aggregate these features. Subsequently, we use a decoder consisting of convolutional layers to reconstruct the fused image. (b): The affine fusion unit is used to fuse the vision signals with the help of the text modality. In our design, the infrared modality is used to generate the weight term $\mu$, while the bias term $\lambda$ is calculated based on the visible images. We expand the spatial dimension of the text features to match the weight term prior to performing the element-wise multiplication operation.
IV-A Network architecture and the affine fusion unit
As shown in Fig. 5 (a), our network is composed of three encoders, i.e., two vision encoders and a text encoder. For the text encoder, we use the popular CLIP model to obtain linguistic features $\psi_{\mathrm{text}}$. On the other hand, the vision encoder part of CLIP contains down-sampling operations, which will impact the image quality when restoring the original size. Thus, to maintain the performance and keep computation efficient, we further design two vision encoders based on the Swin Transformer Blocks [36] for extracting image features $\psi_{\mathrm{ir}}$ and $\psi_{\mathrm{vis}}$. After that, we insert an affine fusion unit into the data processing chain to integrate the extracted features. The final fused image is then reconstructed by a decoder, consisting of convolutional layers.
@@ IV-A Network architecture and the affine fusion unit
Figure 6: Pipeline of the coarse-to-fine association mechanism. CGA and FGA denote the coarse-grained and fined-grained association modules, respectively. The input images together with textual description are first fed into the vision and languange pre-trained model model to obtain the cross-modal heat maps $M_{\mathrm{w}_{i}}$. Then, we aggregate heat maps of the nouns into a single one $\hat{M}$. This heat map will be further used to generate the binary interested map $B$, with subsequent maximizing operation used to yield the initial interested map $\hat{B}$. Finally, we fine-tune this map based on the segmentation techniques and a threshold $\alpha$ in the FGA to produce the final interested map $B_{\mathrm{f}}$.
IV-B Association and training processes
IV-B1 Coarse-to-fine association
As there is no GT image in the fusion task, to control the fused image via linguistic guidance, we need to embed the textual semantics into the loss function design. Thus, we need to generate an association map between the modalities for the training process. However, obtaining a pixel-wise relation map between vision and text modalities can be very expensive [38]. Thus, a coarse-to-fine association mechanism is proposed to alleviate this challenge.
As shown in Fig. 6, aiming to distinguish the user-defined content of interest in the input images from the background, we first construct a coarse-grained association (CGA) module to obtain the heat map $M$ linking each word to visual content, i.e., identifying the degree of correlation between the words and image regions. The heat maps are produced by a vision-and-language pre-trained model. For each noun $\mathrm{w}_{j}$, we fuse their heat maps into a single one $\hat{M}(I)$ by taking the maximum value at each location of these heat maps, i.e.
@@ IV-B1 Coarse-to-fine association
Taking this into consideration, we propose a subsequent fine-grained association (FGA) module to fine-tune the interest map. Such association is determined based on $Q$ instances of source images, e.g., persons, trees and cars, from a segmentation model. We first calculate the overlapping ratio $r_{i}$ between these instances and the interest map, i.e.
$r_{i}=\frac{\sum_{x=1}^{H}\sum_{y=1}^{W}(O_{i}(x,y)\cdot\hat{B}(x,y))}{\sum_{x=1}^{H}\sum_{y=1}^{W}O_{i}(x,y)},$ (3)
where $i\in\{1,2,...,Q\}$ , $H$ and $W$ are the height and width of the input image, respectively.
Figure 7: An illustration of the training process of the proposed model. In our implementation, we assign different levels of salience for the interested and irrelevant regions, based on activity level maps and information measurement, to obtain the weights for the input images. Accordingly, a fusion network is optimized by using a loss function reflecting the salience differences, with the input text information used to guide the fusion process.
Finally, we add the relevant instance segmentation results together to generate the final interest map $B_{\mathrm{f}}$. Here, an instance $O_{i}$ is considered relevant to the interest map if the corresponding overlapping ratio is greater than $\alpha$. Hence, the final interest map is defined as:
@@
A discussion about the threshold $\alpha$ will be presented in Section. V-E2.
@@
IV-B2 Training
As shown in Fig. 7, during the training process, our loss function is designed to focus on the final interest map $B_{\mathrm{f}}$. The total loss is formulated as
@@ IV-B2 Training
where $Loss_{\mathrm{int}}$ and $Loss_{\mathrm{irr}}$ stand for the region of interest and the irrelevant region loss components. As the thermal radiation information can be captured by the pixel intensity distribution, while the texture details in the RGB images are reflected in the gradient information [39, 40, 14], we use such criterion together with the principles of the dominant modality theory in the affine fusion unit to define these two loss functions. Specifically, the weights $w(B_{f},\phi(I))$ and $p(\overline{B_{f}},\nabla\theta(I))$ are adopted in the region of interest and irrelevant region, respectively, to measure the salience degree of source images. $\phi(I)$ and $\theta(I)$ denote the image features obtained from pre-trained models. $\nabla$ is the gradient operator and $\overline{B_{f}}(x,y)=1-B_{f}(x,y)$.
Firstly, in the region of interest, we use the spatial attention strategy adopted in [41, 42] to preserve the salient information from the infrared and visible images. However, different from these methods, we expect to realize the fusion process in an end-to-end manner. To achieve this objective, we adopt the pre-trained autoencoder network [10] to extract the deep image features and use $\phi_{i}(I)$ to denote the $i$-th channel of the feature maps. Subsequently, we conduct the $l_{1}$ normalisation on these features to calculate the activity level map
$A(I)(x,y)=\|\phi_{1:C}(I)(x,y)\|_{1},$ (6)
where A denotes the activity level map and $C$ indicates the number of channels. The activity level map can quantify the salient content of input images [43]. Since information with a higher response value is favoured by this strategy, the region of interest can be well highlighted accordingly. Following this, we conduct a softmax operation on $A(I_{\mathrm{ir}})$ and $A(I_{\mathrm{vis}})$ to obtain the pixel-level weight maps $w_{\mathrm{ir}}$ and $w_{\mathrm{vis}}$.
Next, we do not follow the two-stage autoencoder-based methods. Instead, in the region of interest, we apply these two weights in the loss function to dynamically control the proportion of pixel information transferred from the input images into the fusion result. Thus, we obtain the formulation of this loss function
where $\|\cdot\|_{F}$ is the Frobenius norm.
On the other hand, for the irrelevant regions, the information measurement approach [44] has been shown to handle the visual texture details well. We tend to retain this information in the irrelevant regions as supplementary for the salient interest part. Since we maintain significant contrast for the fused images, such difference in design can also benefit the downstream vision task to better index the position of the foreground objects.
Specifically, we use a similar strategy to control the mixing of the source images in the non-salient areas. Firstly, we estimate the information content of input images using a pre-trained VGG-19 network [45]. This measurement is defined as
where $\theta_{i}^{j}(I)$ denotes the $j$-th channel of the feature maps obtained before the $i$-th max-pooling layer of the VGG model, $n_{i}$ is the number of the feature maps in the corresponding convolutional layer. $DS(X,s)$ denotes the down-sampling operation with the scale factor of $s$ on matrix $X$.
We apply a softmax operation to this information measurement to obtain the mixing weights $p_{{ir}}$ and $p_{{vis}}$. Note that, the information measurement is calculated based on the whole input image. Thus, different from the pixel-level weights in the area of interest, the irrelevant region loss is formulated as follows:
Figure 8: An illustration of the textual attention assessment $Q+$. In this metric, we aim to use the fusion prior and complete description from the volunteers to improve the existing average-based fusion metrics. ($C_{t}$ denotes the confidence value obtained based on the coarse-to-fine association mechanism)
IV-C Textual attention assessments
@@
As shown in Fig. 8, by utilizing the textual modality to describe the region of interest within the input images, we can embed a prior fusion knowledge into the IQAs to evaluate the fusion results. That is, depending on the purpose of the fusion task, we select salient information from the infrared modality and texture details from the visible images in the regions of interest and other areas, respectively, to construct a text-guided fused image $I_{\mathrm{tf}}$. We generate this image in a straightforward manner:
@@
V Experimental results and analysis
V-A Experimental setting
In this paper, our TextFusion is used to implement the infrared and visible image fusion task. As there is no available infrared and visible image dataset with the corresponding text modality, we carefully annotated data from several RGBT datasets, $i.e.$, TNO, RoadScene and LLVIP [46, 44, 47], to build a benchmark text-guided image fusion dataset (IVT). The final database which contains $11,405$ textual annotations will be released to the research community.
During the training process, 10,000 RGBT and text triples from the IVT-LLVIP dataset were used to optimize the fusion model. The segmentation and the vision-and-language models are only pre-trained on the ImageNet classification tasks and the masked language modelling tasks [48, 30]. In the test phase, the entire IVT-TNO and IVT-RoadScene datasets, along with the remaining images from the IVT-LLVIP dataset, are used to evaluate the performance. For the quantitative comparison, some widely used metrics, i.e., an objective image fusion performance measure $Q_{abf}$, structural similarity (SSIM), and visual information fidelity (VIF) [49, 50] are chosen to measure the similarity of the fused image and source input from different perspectives. Besides, we also include two non-reference metrics, i.e., spatial frequency (SF) and standard deviation (SD) to evaluate the image quality of the fusion results. Larger values of these metrics indicate better fusion performance.
The comparison algorithms include the mainstream paradigms that emerged in the last few years, namely a GAN-based method, TarDAL [23], an autoencoder-based method, RFN-Nest [12], a self-evolution paradigm MUFusion [22], two methods combined with other vision tasks ReCoNet [51] and MetaFusion [23], a diffusion-based model DDFM [52] and a novel representation guided network, LRRNet [53]. The experiments are conducted on an NVIDIA GeForce RTX 3090 GPU.
@@ V-B Controllability characteristics
TABLE I: The average detection precision of different fusion methods obtained on images from the IVT-LLVIP dataset. (Bold: best, Bold: second best)
@@
Figure 11: Sample qualitative results of our TextFusion compared with other state-of-the-art methods on the IVT dataset. As shown in the highlighted regions, our method can present significant thermal information while retain clearer textual details form the RGB images, compared with other advanced approaches.
TABLE II: Quantitative results of different image fusion methods obtained on the IVT dataset. ($mRank+$ and $mRank$ denote the mean rankings of the three textual attention and traditional metrics).
@@ V-C Pedestrian detection
Apart from the results visualisation, we also conducted quantitative experiments on this task. As presented in Table I, with or without text input, our TextFusion always ranks first among these advanced image fusion methods.
@@
V-D Image fusion results on the IVT dataset
@@
As shown in Table II, compared to all advanced image fusion methods, our TextFusion delivers the best performance in terms of mRank on these three datasets. This demonstrates that, the results produced by our TextFusion maintain a high correlation with input images, and that our method is able to handle different fusion scenarios robustly. At the same time, as our method produces fused images with higher definition, it always achieves the best or second-best performance in terms of the no-reference metrics SF and SD. This objectively measured performance is consistent with the observations made in the qualitative evaluation.
@@
An analysis of the textual attention metrics: With our textual attention metrics, we aim to better measure the contributions of the source images to the fusion result quality controlled with the help of linguistic semantics. In this section, we combine the qualitative and quantitative experiments to demonstrate the effectiveness of the improved metrics. As shown in Fig. 11, the image fusion results of RFN-Nest suffer from the bias issue, i.e., the output either excessively preserves information from the RGB image or emphasizes prominent thermal radiation content at the expense of losing texture details. However, the traditional image fusion evaluation techniques, by evenly weighting the contributions from the source images, obscure this drawback. As a result, RFN-Nest still manages to achieve the second-best performance in terms of SSIM (Table II). By contrast, our assessment (SSIM+) is able to identify this bias, with its average fusion ranking being suppressed to a certain extent (it drops from 6.00 to 6.67). Meanwhile, as MetaFusion and LRRNet can simultaneously preserve the important content from both modalities (though with lower quality than that produced by our method), their mRank benefits from the proposed assessments. In general, our textual attention metric can objectively identify the methods that fall short in certain aspects and promote those aligning with the fusion task objectives. This ensures that numerical results reflect more accurately the behaviour of image fusion approaches.
@@
TABLE III: Quantitative results of the ablation study relating to the affine fusion unit.
@@
(c) $\mu$-vis $\lambda$-ir Minor 0.491 0.657 0.637 17.897 45.175
Ours ($\mu$-ir $\lambda$-vis) Major 0.527 0.656 0.744 18.872 49.578
V-E Ablation experiments
In this section, we present ablation experiments and visualisation results to highlight the importance of various components in our TextFusion approach.
V-E1 Ablation study relating to the affine fusion unit